If tood is 10 years older than jamal. The sum of their ages is 64. How old is Todd and jumal
Found 2 solutions by josgarithmetic, algebrahouse.com:
Answer by josgarithmetic(39629)
Answer by algebrahouse.com(1659)

j = Jamal's age
j + 10 = Todd's age {Todd is 10 years older than Jamal}

j + j + 10 = 64 {the sum of their ages is 64}
2j + 10 = 64 {combined like terms}
2j = 54 {subtracted 10 from each side}
j = 27 {divided each side by 2}
j + 10 = 37 {substituted 27, in for j, into j + 10}

Jamal is 27
Todd is 37

For more help like this, visit: www.algebrahouse.com